OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how to use the distributive property to simplify the expression below. 5(x + 2) Part 2: Write the simplified expression

OpenStudy (anonymous):

The multiplication of a value; whether it be: a number, variable, or expression, by a group of numbers undergoing a numerical operation within a contained: bracket, parenthesis, or brace(real and implied), can be multiplied to each value contained within. Simplified When multiplying a value by a set of values or expression(s) inside a mathematical container undergoing addition or subtraction, this outside value must be multiplied independently to each value in the numerator contained within. Note: A non-rational expression has an implied denominator of 1. 5(x + 2) --> (5*x +5*2) = 5x+10

OpenStudy (anonymous):

Not to make it more complicated by adding that line about the numerator. Adding this to your answer will have you covered for fractions. 5(x+2) = 5[(x+2)/1] you wouldn't multiple the 5 by the one or anything in the denominator, unless you are multiplying by a fraction. (1/5)[(x+2)/1] = (x+2)/5 Which is equivalent to x/5 + 2/5